OpenStudy (anonymous):

Solve for x. √6x+7+2x=3x A. x=-1 B. x=7,-1 C. x=1 D. x=7

OpenStudy (anonymous):

\[\sqrt{6x}+7+2x=3x\] is this your equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

It's best to split the square roots at the start, \[\sqrt{6}\sqrt{x}+7+2x=3x\] now subtract 2x on both sides, \[\sqrt{6}\sqrt{x}+7=3x-2x\] this would give us 1 x on the right side, now subtract the 7 on both sides, \[\sqrt{6}\sqrt{x}=x-7\] or really you don't have to split the square root but it's fine, now square both sides so you get, \[6x=(x-7)^2 \implies 6x=x^2-14x+49\]

OpenStudy (anonymous):

Now we get rid of the 6x \[(x-10)^2=51 \implies x -10 = \pm \sqrt{51} \implies x = 10\pm \sqrt{51}\]

OpenStudy (anonymous):

And this means none of your solutions are right because you probably gave me the wrong equation

OpenStudy (anonymous):

oh yeah my bad the square root is over the 7 as well @iambatman

OpenStudy (anonymous):

Well try it yourself now, it's basically the same process :)

OpenStudy (anonymous):

See what you get and if you have trouble let me know.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got c

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

i have trouble @iambatman

OpenStudy (anonymous):

Ok tell me how you got it..